# VideoAuteur: Towards Long Narrative Video Generation

### v1

@@ Appendix F Generated Video Examples
Appendix G Limitations
G.1 Noisy “Actions” from ASR
@@ G.2 Failure Cases
Video Generation Model: Unrealistic Hallucination
.Unrealistic hallucination occurs when a video generation model produces content inconsistent with the intended narrative. In Figure 20(a), the action ”placing the fried chicken into an oven set to preheat” is misrepresented as frying chicken in a pan, with an unrealistic increase in the quantity of chicken, showing a lack of object continuity. In Figure 20(b), the action ”adding a drizzle of sauce to a plate of grilled skewers” introduces an illogical appearance of new grilled food items, deviating from the intended action and disrupting narrative coherence.

### v2

@@ Appendix G Generated Video Examples
Appendix H Limitations
H.1 Noisy “Actions” from ASR
@@ H.2 Failure Cases
Video Generation Model: Unrealistic Hallucination.
Unrealistic hallucination occurs when a video generation model produces content inconsistent with the intended narrative. In Figure 23(a), the action ”placing the fried chicken into an oven set to preheat” is misrepresented as frying chicken in a pan, with an unrealistic increase in the quantity of chicken, showing a lack of object continuity. In Figure 23(b), the action ”adding a drizzle of sauce to a plate of grilled skewers” introduces an illogical appearance of new grilled food items, deviating from the intended action and disrupting narrative coherence.